What value(s) of x will make this equationbelow true?
2x - 6 = 3x + 1 - x - 7 ?

Answer by josgarithmetic(39617)

None.
Try solving for x and find what happens.

2x-6=3x+1-x-7
2x-6=(3-1)x+1+7
2x-6=2x+8
and you would want to get all terms of x on one side and all constants on the other side,
2x-6-2x+6=2x-2x+8+6
0=0+14
IMPOSSIBLE.

No solution!